# Infrared spectra of methane-containing ice mixtures for JWST data analysis

### v1

@@ Abstract
Context. Solid methane (CH4) is an important molecule in interstellar and planetary environments, serving as a precursor to complex organic compounds and a potential biosignature in exoplanetary studies. Despite its significance, laboratory data on low-temperature phase of methane below 10 K remain limited.
Aims. We aim to obtain spectra of methane in binary mixtures at 10 K and compare it to the spectra obtained at 6.7 K. These temperatures correspond to phases II and II* of pure methane and are representative of dark molecular clouds and protostars at early stages. We also aim to test the obtained data applicability to JWST data interpretation.
Methods. Laboratory reference spectra were obtained on the ISEAge setup via FTIR spectroscopy in transmission mode. A weighted minimization is used for the fitting.
Results. We present infrared spectra with corresponding band strengths of pure methane and binary mixtures with methane: CH4:H2O, CH4:CO2, CH4:CH3OH, CH4:NH3 at 6.7 K and 10 K showing a 20% increase in mixtures compared to commonly used 10 K band strength value of pure methane. We also test the usability of the spectra on open JWST data by probing the spatial distribution of methane in B335. We also present additional experiments concerning the phase transition of methane between phase II* and phase II.
Conclusions. Our results reveal distinct spectral features for methane in non-H2O environments, enabling more accurate interpretation of JWST observations. The dataset of spectra, publicly available on Zenodo, can be used for fitting JWST data.
@@
Solid methane is the most abundant and one of the simplest hydrocarbon molecules found in prestellar objects and protostars (Öberg et al., 2008). Both theoretical (Hassel et al., 2008; Harada and Herbst, 2008; Carder et al., 2021) and experimental (Baratta et al., 2002; Hodyss et al., 2009; Abplanalp et al., 2018; Zhu et al., 2018) studies consider methane a precursor for more complex carbonaceous molecules. Also, it can possibly be considered as a biosignature for exoplanetary studies (Krissansen-Totton et al., 2018; Thompson et al., 2022; Glass and Hörst, 2025). Methane is a part of the atmosphere of planets, e.g. Jupiter and Mars (Gautier et al., 1982; Krasnopolsky et al., 2004) and exoplanets, e.g. HD 209458b and HR 8799 b (Swain et al., 2009; Barman et al., 2015).
@@
The first discovery of interstellar solid methane occurred by the ground-based infrared echelle spectrograph (IRSHELL) (Lacy et al., 1991). Methane is usually identified on infrared (IR) spectra using 7.7 m (1300 cm-1) deformation mode absorbance band. Boogert et al. (1996) performed the first fit of methane absorption band at 7.7 m with CH4:H2O ice spectrum and proposed the CH4:CH3OH and CH4:NH3 mixtures as alternatives. Based on shape and position of absorbance band it was concluded that methane is contained within the polar layer of ice mantles on interstellar dust (Boogert et al., 1998). Methane was identified in a number of objects in Infrared Space Observatory (ISO) data review (Gibb et al., 2004), where spectra of pure methane and methane in mixture with H2O were used. Later, methane abundance was constrained using Spitzer data, where methane was fitted with CH4:H2O mixtures (Öberg et al., 2008). Currently, research on methane continues with the recent James Webb Space Telescope (JWST) data. The deformation mode of solid methane at 7.7 m is described by methane-bearing H2O ices: at 15 K (H2O:CH4 = 10:1111Mixtures used in fitting procedures in the cited work are shown with this notation, Rocha et al., 2024) in IRAS 2A and IRAS 23385+6053, at 13 K (H2O:CH4 = 100:6, Chen et al., 2024) in B1-c, at 12 K (H2O:CH4 = 10:1, Rocha et al., 2025) in Ced 110 IRS4, H2O and CO2 environment at 12 K222Note that the spectrum that is referred to in fitting procedure is not listed in the original paper (Rocha et al., 2017) (H2O:CO2:CH4 = 10:1:1, McClure et al., 2023) in NIR38 and J110621, using a complex mixture at 10 K (H2O:CH3OH:CO2:CH4 = 0.6:0.7:1:0.1, Yang et al., 2022) in IRAS 15398–3359 and H2O:CH4 = 10:1 at 8 K and CO2:CH4 = 5:1 at 27 K (Nakibov et al., 2025) in IRAS 23385+6053. According to these findings on JWST data the methane abundance in cold clouds and star-forming regions is 1.6–3.7% with respect to solid H2O.
@@
There are three phases of solid methane under vacuum conditions: phase I is stable above 20.4 K, phase II — below 20.4 K (e.g. Gerakines and Hudson, 2015b), and the third phase, stable if methane is deposited below 10 K (Colwell et al., 1962; Krupskii et al., 1973; Emtiaz et al., 2020). Following Emtiaz et al. (2020) we refer to this third phase as crystalline phase II*. However, the literature spectra of methane typically used for interpretation of observational data are obtained at a temperature of 10 K and above, and the laboratory data on ice analogues are present as water-based mixtures for phases I and II (Hudson and Moore, 1999; Ehrenfreund et al., 1999; Rocha et al., 2017; Mifsud et al., 2023). Binary mixtures of methane with H2O, CO, CH3OH, NH3, CO2, N2, O2 for the first time were described in Boogert et al. (1997) at 10 K yet the spectral data needed for fitting is unavailable. There are also reflection-mode studies (e.g. Kaiser et al., 2014; Maity et al., 2014; Förstel et al., 2017; Bergantini et al., 2018) that characterize methane photochemistry using relevant laboratory ices, however, their spectra are not directly comparable to astronomical observations of ices in transmission.
@@
In this paper we present laboratory spectra of binary methane-containing ices in transmission regime, that are currently missing in the literature. These mixtures correspond to the environment of the polar layer of ice on the dust grains and aimed for interpretation of JWST data. The ices have astrochemically relevant component ratios: CH4:H2O = 1:10, CH4:CO2 = 1:5 and 1:15, CH4:CH3OH = 1:3, CH4:NH3 = 1:3. The set of mixtures is obtained at a substrate temperature of 10 K. However, in cold prestellar cores temperature may reach as low as 6 K, which is confirmed by, e.g. observations of gas and dust in centres of isolated cold dense cores L183 and L1544 (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015). Also, physical modeling of L1544 resulted in the dust temperature in the core center equal to 6.5 K (Keto and Caselli, 2010). The presence of dust colder than 10 K is also confirmed by observations obtained with SCUBA at 850 m (Evans et al., 2001; Pattle et al., 2015). Therefore the set was extended with ices grown at 6.7 K with same composition. The extended set explores the difference between the ices grown at 6.7 K and 10 K substrate temperatures, which correspond to crystalline II* and crystalline II phases of pure methane. The transmission spectra of methane in binary mixtures grown at 6.7 K and CH4:CH3OH, CH4:NH3 at 10 K are presented here for the first time. The band strengths of methane in various environment are provided as well.
@@
In our previous paper (Nakibov et al., 2025) we described gaseous and solid methane in IRAS 23385+6053 using some of the spectra presented here. In this paper we show the utility of our new infrared spectra by exploring the spatial distribution of methane in JWST spectra of the variable protostar B335-IRS (hereafter B335). B335 is an isolated nebula at a distance of 90–120 pc with a young class 0 protostar IRAS19347+0727 (age yr, Yıldız et al., 2015). It is associated with protostellar high-velocity Herbig-Haro (HH) jet (Reipurth et al., 1992; Gâlfalk and Olofsson, 2007). This source has a high level of accretion and two outflows (Bjerkeli et al., 2019) and had a luminosity increase by a factor of 5–7 (Evans et al., 2023). This source is a natural laboratory to study ice chemistry at the early stages of star formation. While the sublimation of complex ices and the rich gas-phase chemistry near the heated core was observed at millimeter wavelengths (Lee et al., 2025), we focus here on the solid methane features at 7.7 m located within the outflow cavity wall. This allows us to probe the composition of ices in the cold envelope of a young protostar.
@@ 2 Experimental methods
All experiments on growing analogues of interstellar ice were carried out on Ice Spectroscopy Experimental Aggregate (ISEAge) that utilizes ultra-high vacuum setup. A detailed description of our installation is provided in Ozhiganov et al. (2024). Briefly, the molecules are deposited on the Ge substrate from two sides via ‘background deposition’. The 6 mm diameter circle area (S=0.28 cm2) of the window is available for the deposition of the ice. The substrate is mounted on a cryogenic finger of a closed cycle He cryostat, and the whole chamber is continuously evacuated to the base pressure of × mbar. The operating temperatures are within the 6.7–305 K range with an error 0.1 K. The setup is equipped with a Stanford Research Systems RGA200 quadrupole mass spectrometer and Thermo Scientific Nicolet iS50 Fourier Transform Infrared (FTIR) spectrometer operating within a 4000–630 cm-1 (2.5–15.9 m) with 1 cm-1 resolution. The points in the collected spectra are spaced at 0.12 cm-1 intervals due to the Zero Filling performed on the interferometer level. The infrared spectra are collected in transmission regime. The FTIR spectrometer and optical path outside the main chamber are flushed by pure N2. Gaseous species are introduced into the main chamber via two independent leak valves.
@@
We followed the same experimental procedure for growing of all ices. The component ratio was controlled during the deposition by means of the quadrupole mass spectrometer (QMS). In all experiments both components of a binary mixture were deposited simultaneously. The first component introduced into the main chamber was the pure methane and the second component was a matrix molecule introduced shortly after the desired ion current of methane was reached. The mixture deposition time started when ion currents for both components reached the desired values. This approach allows to correct the influence of ion interference and overlapping ion currents, and maintain the same ion current for methane across all the experiments. The ion current values were taken from the calibration curves obtained for our setup via the procedure similar to Slavicinska et al. (2023). Briefly, the rate at which gaseous molecules are adsorbed on the substrate depends nonlinearly on the ion current and is unique to each molecule. By growing pure ices at different ion currents we can obtain this dependency via experiment allowing for a more precise control over the components ratio of mixed ices. The rates for the calibration curves are determined by measuring the absorbance on the IR spectrum output using the literature band strength values. Since the depositions were conducted on the Ge substrate cooled to 10 K and 6.7 K we used the band strength for methane in phase II (corrected for density) from Bouilloud et al. (2015) and for methane in phase II* from Gerakines and Hudson (2015b), correspondingly.
@@
For both sets of experiments the methane deposition rate was fixed at cm-2 s-1 and the second component deposition rate varied based on the desired components ratio. Deposition continued for 120 minutes. Thus, we get a column density of cm-2 for CH4. This approach enables the comparison of methane deformation (1300 cm-1) and stretching (3000 cm-1) modes in various molecular environment. The IR spectra were recorded every 45 seconds (an average of 32 scans). For the analysis in each experiment the last recorded spectrum of deposition was used to avoid the secondary adsorption effect after closing the leak valve. This effect is present when working with volatile species such as CH4, CO or N2 both in static conditions and during the warm up, see Appendix A for details. For calculations of peak positions and areas spectra are baseline-corrected for overall slope and local thin-film interference artifacts. Repeated measurements (50 iterations) with randomized local baseline adjustments are analyzed to minimize noise effects. The final values are derived from the averaged results.
@@
The compounds used are as follows: CH4 (99.999 %, Ugra-PGS), CO2 (99.9999 %, Ugra-PGS), NH3 (99.9 %, UralKrioGaz), deionized H2O, CH3OH (99.8 %, Vekton). Gaseous CO2 and CH4 are introduced directly into the dosing lines from the commercially acquired gas bottles. H2O, NH3 and CH3OH were purified by a freeze-pump-thaw cycle three times before each experiment. Before each deposition every compound is checked for purity by means of mass-spectrometry.
@@ 3.1 Binary mixtures with CH4
In this section we present a series of experiments on obtaining the binary ices of methane with molecules that constitute the polar layer of the interstellar ice mantles. We obtained binary mixtures at a temperature of 10 K, with methane molecules embedded in a matrix of H2O, CO2, CH3OH or NH3. The component ratio was chosen to qualitatively reflect current astrochemical view on species abundances in interstellar ices in cold dense molecular clouds (Boogert et al., 2015; Goto et al., 2021; McClure et al., 2023; Jiménez-Serra et al., 2025). Additionally, the comparative set was obtained at the minimum stable temperature of the ISEAge setup at 6.7 K. Methane absorption bands in mixtures at 6.7 K spectra differ both from pure methane and 10 K mixtures features. We present the spectra of mixtures for both temperatures in Fig. 1. Additional experiments concerning pure methane and its transition between phases II* and II are presented in Appendix A.
@@
The methane absorption bands change shape and position in mixed ices, the components ratio and peak position of 3000 cm-1 stretching and 1300 cm-1 deformation mode are presented in Table 1. The band strengths for the deformation mode of methane in various mixtures are also included. Methane deformation mode in CH4:CO2 = 1:5 mixture stands out because of its structure. It displays an unresolved double peak both in 6.7 K and 10 K experiments, which becomes resolved on heating after 30–40 K (Boogert et al., 1997; Nakibov et al., 2025). Moreover, this mixture displays the most distinction between the two temperatures in both vibrational modes. This is caused by pure CO2 having an amorphous and crystalline form as shown in Gerakines and Hudson (2015a). For our fixed column density of 2.121017 cm-2 pure CO2 fully crystallizes at 10 K and only partially at 6.7 K. An addition of methane, however, prevents CO2 from crystallizing at 6.7 K. To further investigate this we obtained more diluted mixtures: CH4:CO2 = 1:15 at 6.7 K and 10 K, the differences persisted despite the higher CO2 concentration. We report, therefore, the peak shape of methane differs significantly in crystalline and amorphous CO2. We also note that the double peak gets resolved at 10 K in CH4:CO2 = 1:15 mixture, which is in line with Hudgins et al. (1993), where the double structure is observed in CH4:CO2 = 1:20 mixture at 10 K.
@@
Methane mixtures with CH3OH and NH3 show similar features in 1300 cm-1 region and have close band strengths values. The features are redshifted by 2 cm-1 from the H2O:CH4 feature. This opens a possibility of singling out methane not related to water ice in observational spectra, an idea first introduced by Boogert et al. (1996). This is important, as normally methane is assumed to be formed simultaneously with water and ammonia ice via hydrogenation of carbon atoms at the onset of the formation of icy grain mantles (Qasim et al., 2020; Lamberts et al., 2022; Ferrero et al., 2023). However, methane can also be formed as a product of photoprocessing or ion-bombardment of methanol-rich ices (Islam et al., 2014) at later stages of ice evolution or via electron-induced radiolysis (Sullivan et al., 2016). The presented data allows potentially to explore the relative importance of both scenarios of methane formation.
@@
| CH4 | - | 8 | 3010.0 | 1296.9 | 1.04×10-17aaaaThe band strength is taken from Gerakines and Hudson (2015b); |

### v2

# Infrared spectra of methane-containing ice mixtures for JWST data analyses

@@ Abstract
Context. Solid methane (CH4) is an important molecule in interstellar and planetary environments, serving as a precursor to complex organic compounds, and it is a potential biosignature in exoplanetary studies. Despite its significance, laboratory data on the low-temperature phase of methane below 10 K remain limited.
Aims. We obtained spectra of methane in binary mixtures at 10 K and compared them to the spectra obtained at 6.7 K. These temperatures correspond to phases II and II* of pure methane and are representative of dark molecular clouds and protostars in early stages. We also tested whether the data we obtained can be applied to interpret JWST data.
Methods. Laboratory reference spectra were obtained with the ISEAge setup via Fourier transform infrared spectroscopy in transmission mode. A weighted minimization was used for the fitting.
Results. We present infrared spectra with corresponding band strengths of pure methane and binary mixtures with methane: CH4:H2O, CH4:CO2, CH4:CH3OH, and CH4:NH3 at 6.7 K and 10 K. They show an increase of 20% in mixtures compared to the commonly used 10 K band strength value of pure methane. We also tested whether the spectra can be used on open JWST data by probing the spatial distribution of methane in B335. We also present additional experiments concerning the phase transition of methane between phase II* and phase II.
Conclusions. Our results reveal distinct spectral features for methane in non-H2O environments that enable a more accurate interpretation of JWST observations. The dataset of spectra is publicly available on Zenodo and can be used for fitting JWST data.
@@
Solid methane is the most abundant and one of the simplest hydrocarbon molecules found in pre-stellar objects and protostars (Öberg et al., 2008). Both theoretical (Hassel et al., 2008; Harada and Herbst, 2008; Carder et al., 2021) and experimental (Baratta et al., 2002; Hodyss et al., 2009; Abplanalp et al., 2018; Zhu et al., 2018) studies have considered methane a precursor for more complex carbonaceous molecules. It is also considered a biosignature for exoplanetary studies (Krissansen-Totton et al., 2018; Thompson et al., 2022; Glass and Hörst, 2025). Methane is a part of the atmosphere of planets, for instance Jupiter and Mars (Gautier et al., 1982; Krasnopolsky et al., 2004), and exoplanets, for example HD 209458b and HR 8799 b (Swain et al., 2009; Barman et al., 2015).
@@
The first discovery of interstellar solid methane was made with the ground-based infrared (IR) echelle spectrograph IRSHELL (Lacy et al. 1991). Methane is usually identified on IR spectra using the 7.7 m (1300 cm-1) deformation mode absorbance band. Boogert et al. (1996) performed the first fit of the methane absorption band at 7.7 m with the CH4:H2O ice spectrum and proposed the CH4:CH3OH and CH4:NH3 mixtures as alternatives. Based on the shape and position of the absorbance band, it was concluded that methane is contained within the polar layer of ice mantles on interstellar dust (Boogert et al., 1998). Methane was identified in a number of objects in a review of Infrared Space Observatory data (Gibb et al., 2004), where spectra of pure methane and methane mixed with H2O were used. Later, the methane abundance was constrained using Spitzer data, where methane was fitted with CH4:H2O mixtures (Öberg et al., 2008). Currently, research on methane continues with the recent James Webb Space Telescope (JWST) data. The deformation mode of solid methane at 7.7 m is described by methane-bearing H2O ices: at 15 K (H2O:CH4 = 10:1111Mixtures used in fitting procedures in the cited work are shown with this notation.; Rocha et al., 2024) in IRAS 2A and IRAS 23385+6053, at 13 K (H2O:CH4 = 100:6; Chen et al., 2024) in B1-c, at 12 K (H2O:CH4 = 10:1; Rocha et al., 2025) in Ced 110 IRS4, for an H2O and CO2 environment at 12 K222Note that the spectrum that is referred to in the fitting procedure is not listed in the original paper (Rocha et al., 2017). (H2O:CO2:CH4 = 10:1:1; McClure et al., 2023) in NIR38 and J110621, using a complex mixture at 10 K (H2O:CH3OH:CO2:CH4 = 0.6:0.7:1:0.1; Yang et al., 2022) in IRAS 15398–3359, and H2O:CH4 = 10:1 at 8 K and CO2:CH4 = 5:1 at 27 K (Nakibov et al., 2025) in IRAS 23385+6053. According to these findings, the methane abundance in cold clouds and star-forming regions is 1.6–3.7% with respect to solid H2O.
@@
There are three phases of solid methane under vacuum conditions: phase I is stable above 20.4 K, phase II is stable below 20.4 K (e.g., Gerakines and Hudson, 2015b), and the third phase is stable when methane is deposited below 10 K (Colwell et al., 1962; Krupskii et al., 1973; Emtiaz et al., 2020). Following Emtiaz et al. (2020), we refer to this third phase as the crystalline phase II*. However, the literature spectra of methane that were typically used to interpret observational data were obtained at temperatures of 10 K and above, and the laboratory data on ice analogs are presented as water-based mixtures for phases I and II (Hudson and Moore, 1999; Ehrenfreund et al., 1999; Rocha et al., 2017; Mifsud et al., 2023). Binary mixtures of methane with H2O, CO, CH3OH, NH3, CO2, and N2, O2 were described for the first time by Boogert et al. (1997) at 10 K, but the spectral data needed for fitting are not available. There are also reflection-mode studies (e.g., Kaiser et al., 2014; Maity et al., 2014; Förstel et al., 2017; Bergantini et al., 2018) that characterized methane photochemistry using relevant laboratory ices, but their spectra are not directly comparable to astronomical observations of ices in transmission.
@@
We present laboratory spectra of binary methane-containing ices in the transmission regime that are currently missing in the literature. These mixtures correspond to the environment of the polar layer of ice on the dust grains and are meant to help us interpret JWST data. The ices have astrochemically relevant component ratios: CH4:H2O = 1:10, CH4:CO2 = 1:5 and 1:15, CH4:CH3OH = 1:3, and CH4:NH3 = 1:3. The set of mixtures was obtained at a substrate temperature of 10 K. However, in cold pre-stellar cores, temperatures can reach as low as 6 K, which was confirmed, for instance, by observations of gas and dust in the centers of the isolated cold dense cores L183 and L1544 (Pagani et al., 2004, 2007; Crapsi et al., 2007; Pagani et al., 2015). Moreover, the physical modeling of L1544 yielded a dust temperature in the core center equal to 6.5 K (Keto and Caselli, 2010). The presence of dust colder than 10 K was also confirmed by observations obtained with Submillimetre Common-User Bolometer Array (SCUBA) at 850 m (Evans et al., 2001; Pattle et al., 2015). Therefore, the set was extended with ices grown at 6.7 K with the same composition. The extended set explored the difference between the ices grown at 6.7 K and 10 K substrate temperatures, which correspond to the crystalline II* and crystalline II phases of pure methane. The transmission spectra of methane in binary mixtures grown at 6.7 K and CH4:CH3OH and CH4:NH3 at 10 K are presented here for the first time. The band strengths of methane in various environments are provided as well.
@@
In our previous paper (Nakibov et al., 2025), we described gaseous and solid methane in IRAS 23385+6053 using some of the spectra presented here. In this paper, we show the utility of our new IR spectra by exploring the spatial distribution of methane in JWST spectra of the variable protostar B335-IRS (hereafter B335). B335 is an isolated nebula at a distance of 90–120 pc with a young class 0 protostar, IRAS19347+0727 (age yr; Yıldız et al., 2015). It is associated with a protostellar high-velocity Herbig-Haro jet (Reipurth et al., 1992; Gâlfalk and Olofsson, 2007). This source has a high level of accretion and two outflows (Bjerkeli et al., 2019), and its luminosity increased by a factor of 5–7 on Modified Julian Date (MJD) 56948 (Evans et al., 2023). This source is a natural laboratory for studying ice chemistry in the early stages of star formation. While the sublimation of complex ices and the rich gas-phase chemistry near the heated core were observed at millimeter wavelengths (Lee et al., 2025), we focus here on the solid methane features at 7.7 m that are located within the outflow cavity wall. This allows us to probe the composition of ices in the cold envelope of a young protostar.
@@ 2 Experimental methods
All experiments on growing analogs of interstellar ice were carried out on the Ice Spectroscopy Experimental Aggregate (ISEAge), which uses an ultrahigh vacuum setup. A detailed description of our installation is provided in Ozhiganov et al. (2024). Briefly, the molecules are deposited on the Ge substrate from two sides via background deposition. A circular area of the window with a diameter of 6 mm (S=0.28 cm2) is available for the deposition of the ice. The substrate is mounted on a cryogenic finger of a closed-cycle He cryostat, and the whole chamber is continuously evacuated to the base pressure of × mbar. The operating temperatures are within the 6.7–305 K range with an error of 0.1 K. The setup is equipped with a Stanford Research Systems RGA200 quadrupole mass spectrometer and a Thermo Scientific Nicolet iS50 Fourier transform infrared (FTIR) spectrometer operating within 4000–630 cm-1 (2.5–15.9 m) with 1 cm-1 resolution. The points in the collected spectra are spaced at 0.12 cm-1 intervals due to the zero filling performed on the interferometer level. The IR spectra are collected in the transmission regime. The FTIR spectrometer and optical path outside the main chamber are flushed by pure N2. Gaseous species are introduced into the main chamber via two independent leak valves.
@@
We followed the same experimental procedure to grow all ices. The component ratio was controlled during the deposition by means of the quadrupole mass spectrometer (QMS). In all experiments, the two components of a binary mixture were deposited simultaneously. The first component introduced into the main chamber was the pure methane, and the second component was a matrix molecule introduced shortly after the desired ion current of methane was reached. The mixture deposition time started when ion currents for both components reached the desired values. This approach allowed us to correct for the effect of ion interference and overlapping ion currents and to maintain the same ion current for methane in all the experiments. The ion current values were taken from the calibration curves obtained for our setup in a procedure similar to that described by Slavicinska et al. (2023). Briefly, the rate at which gaseous molecules are adsorbed on the substrate depends nonlinearly on the ion current and is unique to each molecule. By growing pure ices at different ion currents, we obtain this dependence via an experiment, which gives us more precise control over the component ratio of mixed ices. The rates for the calibration curves were determined by measuring the absorbance on the IR spectrum output using the literature band strength values. Since the depositions were conducted on the Ge substrate cooled to 10 K and 6.7 K, we used the band strength for methane in phase II (corrected for density) from Bouilloud et al. (2015) and for methane in phase II* from Gerakines and Hudson (2015b).
@@
For both sets of experiments, the methane deposition rate was fixed at cm-2 s-1, and the second component deposition rate varied based on the desired ratio of the components. The deposition continued for 120 minutes. We obtained a column density of cm-2 for CH4. This approach enabled us to compare the methane deformation (1300 cm-1) and stretching (3000 cm-1) modes in various molecular environments. The IR spectra were recorded every 45 seconds (an average of 32 scans). For the analysis, we used the last recorded deposition spectrum in each experiment to avoid the secondary adsorption effect after closing the leak valve. This effect is present when working with volatile species such as CH4, CO, and N2 in both static conditions and during the warm-up (see Appendix A for details). To calculate the peak positions and areas, the spectra were baseline-corrected for the overall slope and local thin-film interference artifacts. Repeated measurements (50 iterations) with randomized local baseline adjustments were analyzed to minimize noise effects. The final values were derived from the averaged results.
@@
The compounds we used were CH4 (99.999 %, Ugra-PGS), CO2 (99.9999 %, Ugra-PGS), NH3 (99.9 %, UralKrioGaz), and deionized H2O and CH3OH (99.8 %, Vekton). Gaseous CO2 and CH4 were introduced directly into the dosing lines from the commercially acquired gas bottles. H2O, NH3, and CH3OH were purified via a freeze-pump-thaw cycle three times before each experiment. Before each deposition, every compound was checked for purity by means of mass spectrometry.
@@ 3.1 Binary mixtures with CH4
In this section we present a series of experiments we conducted to obtain the binary ices of methane with molecules that constitute the polar layer of the interstellar ice mantles. We obtained binary mixtures at a temperature of 10 K, with methane molecules embedded in a matrix of H2O, CO2, CH3OH, or NH3. The component ratio was chosen to qualitatively reflect the current astrochemical view on species abundances in interstellar ices in cold dense molecular clouds (Boogert et al., 2015; Goto et al., 2021; McClure et al., 2023; Jiménez-Serra et al., 2025). Additionally, the comparative set was obtained at the minimum stable temperature of the ISEAge setup at 6.7 K. Methane absorption bands in mixtures at 6.7 K spectra differ from both pure methane and 10 K mixture features. We present the spectra of the mixtures for both temperatures in Fig. 1. Additional experiments concerning pure methane and its transition between phases II* and II are presented in Appendix A.
@@
The methane absorption bands change shape and position in mixed ices. The component ratio and peak position of the 3000 cm-1 stretching and 1300 cm-1 deformation mode are presented in Table 1. The band strengths for the deformation mode of methane in various mixtures are also included. The methane deformation mode in the CH4:CO2 = 1:5 mixture stands out because of its structure. It displays an unresolved double peak in both the 6.7 K and 10 K experiments, which becomes resolved by heating after 30–40 K (Boogert et al., 1997; Nakibov et al., 2025). Moreover, this mixture displays the strongest distinction between the two temperatures in both vibrational modes. This is caused by the amorphous and crystalline form of pure CO2, as shown by Gerakines and Hudson (2015a). For our fixed column density of 2.121017 cm-2, pure CO2 fully crystallizes at 10 K and crystallizes only partially at 6.7 K. An addition of methane, however, prevents CO2 from crystallizing at 6.7 K. To further investigate this, we obtained more diluted mixtures: CH4:CO2 = 1:15 at 6.7 K and 10 K. The differences persisted despite the higher CO2 concentration. We therefore report that the peak shape of methane differs significantly in crystalline and amorphous CO2. We also note that the double peak becomes resolved at 10 K in the CH4:CO2 = 1:15 mixture, which is in line with the results of Hudgins et al. (1993), who observed the double structure in a CH4:CO2 = 1:20 mixture at 10 K.
@@
Methane mixtures with CH3OH and NH3 show similar features in the 1300 cm-1 region and have similar band strength values. The features are redshifted by 2 cm-1 from the H2O:CH4 feature. This opens the possibility of singling out methane that is not related to water ice in observational spectra, an idea first introduced by Boogert et al. (1996). This is important, as normally methane is assumed to be formed simultaneously with water and ammonia ice via the hydrogenation of carbon atoms at the onset of the formation of icy grain mantles (Qasim et al., 2020; Lamberts et al., 2022; Ferrero et al., 2023). However, methane can also be formed as a product of the photo-processing or ion-bombardment of methanol-rich ices (Islam et al., 2014) in later stages of the ice evolution or via electron-induced radiolysis (Sullivan et al., 2016). The presented data could potentially be used to explore the relative importance of the two scenarios of methane formation.
@@
| CH4 | - | 8 | 3010.0 | 1296.9 | 1.04×10-17aaaaThe band strength is taken from Gerakines and Hudson (2015b). |